OpenStudy (anonymous):

Which of the following is equivalent to log [125] 25 ? 1.076 3/2 2/3 2/5

OpenStudy (mertsj):

Is 125 the base?

OpenStudy (lgbasallote):

quick question...have you tried the calculator?

OpenStudy (zepp):

Do you know the change of base formula?

OpenStudy (mertsj):

\[\log_{125}25=x \] Put it in exponential form.

OpenStudy (mertsj):

Then rewrite 125 and 25 as powers of 5.

OpenStudy (anonymous):

can cheat with change of base. elsewise since \[5^2=25\] and \[5^3=125\] \[125^{\frac{2}{3}}=\sqrt[3]{125}^2=5^2=25\] so your answer is \[\frac{2}{3}\]

OpenStudy (anonymous):

Thanks:)

OpenStudy (lgbasallote):

if you're gonna cheat then might as well do \[\log 25 \div \log 125\] @satellite73

OpenStudy (zepp):

Must read @lgbasallote's tutorial

OpenStudy (anonymous):

zacty what i meant. got a calculator, all you need is change of base and be done

OpenStudy (anonymous):

course your calculator will give you something like \(.666667\)

OpenStudy (anonymous):

i know how to do it, i just dont want to... And i dont mean to be rude but if you dont want to help why comment?

OpenStudy (zepp):

@satellite73 Math -> Frac! :D http://puu.sh/uN3R